A gas stove operates by burning hydrocarbons such as methane (CH4) through the following reaction:

  CH4 (g) + 2 O2 (g) ⟶ CO2 (g) + 2 H2O (g)    △H = -890.8 kJ/mol

If the methane is stored in a container at STP, how much L of methane is necessary to heat 6.76 L of water from room temperature (22.0 °C) to boiling (100.0 °C)? (Hint: the density of water is 1.00 g/mL and the specific heat capacity is 4.184 J/g°C)



1
Expert's answer
2021-10-24T01:25:05-0400

"Q= m\u00d7c\u00d7dT"

-890.8=m× 4.184×78

m= 2.73moles

Moles of methane reacted= 2.73/2= 1.36

Mass= 16×1.36= 21.76g

Volume= 21.76cm3

Volume= 0.02176L